OpenStudy (anonymous):

The angle of elevation of a ladder leaning against a wall is 45°. If the distance from the base of the ladder to the wall is 15 feet, find the length of the ladder. A. 15 feet B. 21.2 feet C. 30.5 feet D. 42.4 feet I think its C?

OpenStudy (hba):

@Anyssah Why do you think it's C ?

OpenStudy (anonymous):

B.\[\sqrt{2*15^2}=15 \sqrt{2}=21.2\text{ feet}\]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

Thank you for not giving me a medal for my brilliant answer.